Model equation for contaminant concentration in a space and apply to the following. A person breathes out CO2 at the rate of 0.3L/min. The concentration of CO2 in the incoming ventilation air is 400 ppm. It is desired to hold the concentration in the room below 1000ppm. Assuming that the air in the room is perfectly mixed, what is the minimum rate of air flow is required to maintain the desired level as per ASHRAE standard 62-1999?. If the same room is designed to occupy 6 persons determine the total amount of CO2 per second breath out and amount of incoming air to meet the ASHRAE standard 62-1999 desired comfortable ambient.
1
Expert's answer
2020-04-19T07:39:55-0400

Person breathes out at the rate 0.3 L/min = "\frac{0.3 L}{60s}= 0.005 L/s" for 6 person "CO_2" expels


out = "6 \times 0.005=0.030 \frac{L}{s}"


The concentration of CO2 is 400 ppm and it has to maintain upto 1000 ppm.


AT STP volume of CO2 is 22.4 L which is equal to 1 mole = 44 g


so mass of CO2 expells out = "0.03\times 44= 1.32 g" /s


400ppm means 400 g of CO2 present in 1000000 particles of air


concentration of incoming air = "400 \times \frac{44}{22.4} = 98.3 \frac{g}{L}"


six person breaths out = "\frac{0.3}{60}\times6\times \frac{44}{22.4}=0.059 g/s"


s per the standards


air flows= "\frac{35.3+98.3Q}{C_o}"